Consider f(x)=x^2+99x+18. find the value of x so that f(2x)=2f(x).
Answer by ewatrrr(24785)

 
Hi,
finding the value of x so that f(2x)=2f(x).
f(x)=x^2+99x+18
(2x)^2 + 99(2x) + 18 = 2x^2 + 198x + 36
4x^2 + 198x + 18 = 2x^2 + 198x + 36
2x^2 - 18 = 0
x^2 = 9
x = ± 3